### v1

@@ Abstract
Predictive tasks on relational databases are critical in real-world applications spanning e-commerce, healthcare, and social media. To address these tasks effectively, Relational Deep Learning (RDL) encodes relational data as graphs, enabling Graph Neural Networks (GNNs) to exploit relational structures for improved predictions. However, existing heterogeneous GNNs often overlook the intrinsic structural properties of relational databases, leading to modeling inefficiencies. Here we introduce RelGNN, a novel GNN framework specifically designed to capture the unique characteristics of relational databases. At the core of our approach is the introduction of atomic routes, which are sequences of nodes forming high-order tripartite structures. Building upon these atomic routes, RelGNN designs new composite message passing mechanisms between heterogeneous nodes, allowing direct single-hop interactions between them. This approach avoids redundant aggregations and mitigates information entanglement, ultimately leading to more efficient and accurate predictive modeling. RelGNN is evaluated on 30 diverse real-world tasks from RelBench Fey et al. (2024), and consistently achieves state-of-the-art accuracy with up to 25% improvement.
@@ 1 Introduction
Predictive modeling over relational data (multiple tables connected via primary-foreign key relations) is central to numerous real-world applications: e-commerce platforms forecast product demand, music streaming services personalize recommendations, and financial institutions assess credit risk. The common strategy to tackle these predictive tasks involves classical tabular machine learning approaches (Chen & Guestrin, 2016) that often require flattening relational data into a single table through manual feature engineering (Kaggle, 2022). This approach is not only labor-intensive but also leads to a substantial loss of predictive signal, as it oversimplifies the interconnected structure of relational data during the flattening process.
@@
To overcome these limitations, Fey et al. (2024) introduced Relational Deep Learning (RDL), a framework that enables end-to-end trainable neural networks to perform predictive modeling directly on relational databases by leveraging their inherent structure. In RDL, relational data is represented as a graph, where each entity is represented as a node, and the primary-foreign key links between entities define the edges. This graph-based representation allows Graph Neural Networks (GNNs)  (Gilmer et al., 2017; Hamilton et al., 2017) to serve as predictive models, capturing complex relational dependencies that traditional methods overlook. Complementing this advancement, RelBench (Robinson et al., 2024) provides the first comprehensive benchmark for evaluating and developing RDL models. By addressing the shortcomings of tabular approaches, RDL not only enhances predictive performance but also establishes a new paradigm in machine learning—one where end-to-end learning frameworks are applied to solve predictive problems on relational data.
Designing effective GNN models for RDL is essential for tackling predictive tasks and supporting critical real-world applications. Relational graphs are large-scale heterogeneous networks whose structure evolves dynamically over time. Existing approaches typically apply standard heterogeneous GNNs directly to these graphs Robinson et al. (2024). However, this can be suboptimal, as these models fail to fully capture the unique structural and relational properties inherent in relational databases. Standard heterogeneous GNNs Schlichtkrull et al. (2018); Hu et al. (2020) are designed for general heterogeneous graphs, where edge types encode direct semantic interactions between entities. In contrast, relational data graphs are structured around primary-foreign key relationships, which define the connectivity between tables rather than semantic meaning. This fundamental distinction significantly affects how information propagates, highlighting the need for models specifically tailored to the structural characteristics of relational databases.
@@
Figure 1: An illustration of the key concepts in our method. (a) The primary-foreign key relation of the rel-f1 dataset. Arrows point from a node with a foreign key to the node with the corresponding primary key. Nodes with zero or one foreign key are marked in blue, and the corresponding foreign key is illustrated by a solid line. Nodes with two or more foreign keys are marked in purple, and the corresponding foreign keys are illustrated by dotted lines. (b) An example of the bridge structure, where standings node is a bridge node. (c) An example of the hub structure, where results node is a hub node. (d) Examples of three atomic routes, where the nodes within each box constitute a distinct atomic route.
@@
Our RelGNN introduces the notion of atomic routes as the foundation of its graph attention mechanism. An atomic route is a sequence of node-types that jointly form a complete information exchange (cf. Fig. 1(d)). We distinguish two scenarios: (i) when a table has a single foreign key referencing another table, and (ii) when a table has multiple foreign keys, acting as a bridge or hub node. In the first case, an atomic route consists of just two node-types and the edge connecting them, whereas in the second case, it comprises a single-hop path grouping the source node, the bridge (or hub) node, and the destination node, thereby capturing all relevant interactions. Although reminiscent of meta-paths in conventional heterogeneous graphs, atomic routes differ fundamentally in their construction: while meta-paths typically rely on domain expert knowledge, atomic routes are systematically derived from primary–foreign key relationships in relational data. Building upon these routes, RelGNN designs composite message passing and attention mechanisms to avoid aggregating irrelevant information, thereby preventing noise accumulation and oversmoothing. Moreover, unlike standard heterogeneous GNNs that require multiple hops for complete information exchange, RelGNN enables direct, single-step communication between the relevant node-types, efficiently extracting critical predictive signals.
We assess the performance of the proposed RelGNN across all tasks in RelBench (Fey et al., 2024), a benchmark which spans seven diverse relational databases covering e-commerce, social networks, sports, and medical platforms. RelBench features 30 real-world predictive tasks cast as entity classification, entity regression, and recommendation. RelGNN surpasses all baselines on 27 of the 30 tasks while performing comparably on the remaining three. Notably, RelGNN achieves more than a 4% improvement over a standard heterogeneous GNN in 15 out of 30 tasks, and provides up to a 25% improvement on the site-success regression task in the rel-trial database.
@@ 2.1 Relational Database
A relational database $(\mathcal{T},\mathcal{L})$ consists of a set of tables $\mathcal{T}=\{T_{1},\ldots,T_{n}\}$ and a set of links between them $\mathcal{L}\subseteq\mathcal{T}\times\mathcal{T}$. Each table is a set $T=\{v_{1},...,v_{n_{T}}\}$, where the elements $v_{i}\in T$ are called rows or entities. Each entity $v\in T$ has a unique primary key $p_{v}$ that distinguishes it from other entities within the table. An entity may also have one or more foreign keys, where each foreign key $\mathcal{K}_{v}\subseteq\{p_{v^{\prime}}:v^{\prime}\in T^{\prime}\text{ and }(T,T^{\prime})\in\mathcal{L}\}$ defines a link between element $v\in T$ to elements $v^{\prime}\in T^{\prime}$, where $p_{v^{\prime}}$ is the primary key of an entity $v^{\prime}$ in table $T^{\prime}$. Besides, an entity may have several attributes $x_{v}$, which represent the informational content of the entity, and an optional timestamp $t_{v}$, indicating when an event occurred. Both primary keys and foreign keys, as well as attributes and the optional timestamp, are columns of the table. A link $L=(T_{\rm fkey}$, $T_{\rm pkey})$ between tables exists if a foreign key column in $T_{\rm fkey}$ references a primary key column of $T_{\rm pkey}$.
For example, in Figue 2(a), the Transactions table has a primary key (TransactionID), two foreign keys (ProductID and CustomerID), one attribute (Price), and a timestamp (Timestamp). Similarly, the Products table has a primary key (ProductID), no foreign keys, three attributes (Description, Image and Size), and no timestamp. The links between tables are illustrated by black connecting lines.
2.2 Relational Deep Learning
Fey et al. (2024) proposed Relational Deep Learning (RDL), a framework that enables end-to-end trainable neural network models to tackle predictive tasks on relational databases. In RDL, relational tables are represented as a temporal, heterogeneous graph, where each table corresponds to a node-type, each entity (row) corresponds to a node, and primary-foreign key links define the edges (See Figure 2 b, c). The information carried by each entity, specified by the columns of the table, is extracted as the initial embedding for the corresponding node. This representation preserves all the information and predictive signals in relational data.
It is important to note that relational data evolves over time as events occur. This temporal aspect is captured by the (optional) timestamp $t_{v}$ attached to each entity $v$. For instance, each transaction in the Transactions table has a timestamp. Many predictive tasks involve forecasting future events, such as predicting the total sales of a product in the next week. Consequently, it is crucial to treat time as an important component in RDL. Specifically, temporal neighbor sampling (Hamilton et al., 2017; Fey et al., 2024) is employed to construct subgraphs around entity nodes at specific seed times determined by the corresponding tasks, with nodes from future timestamps excluded during the sampling to prevent information leakage from future events during training. After that, GNNs can be trained end-to-end on these temporally sampled subgraphs, eliminating the need for manual feature engineering (See Figure 2 d).
2.3 Meta-path in Heterogeneous Graphs
Meta-path  (Sun et al., 2011) is widely used in heterogeneous graphs to capture semantic relationships between different types of entities  (Shang et al., 2016; Dong et al., 2017; Hu et al., 2018; Shi et al., 2018; Wang et al., 2019b; Fu et al., 2020). A meta-path is a sequence of node and edge types in a heterogeneous graph, often designed using expert knowledge to capture meaningful relational patterns. For example, in an academic graph, a meta-path can be defined as ”Author-Paper-Author” to model co-authorship relations, or ”Author-Paper-Conference-Paper-Author” to capture co-conference relations. Despite its widespread use, the meta-path has notable limitations  (Shi et al., 2016; Hu et al., 2020; Shi, 2022). It requires manual selection, which depends on domain expertise and can be biased, leading to suboptimal performance if an inappropriate meta-path is chosen. Additionally, meta-path lacks flexibility, as it cannot easily adapt to changes in graph structure or new relationships. Moreover, designing an effective set of meta-paths for complex graphs can be time-consuming and may fail to capture all relevant interactions, limiting its expressiveness.
3 Method
3.1 Challenges in Message Passing for Relational Deep Learning
Message passing in heterogeneous graphs is designed to aggregate information across nodes and edges of different types. However, relational data graphs exhibit distinct structural properties that set them apart from general heterogeneous graphs. Specifically, in a heterogeneous graph, the fundamental unit of structure is a relation type, typically represented as a triplet of the form (node-type 1, edge-type, node-type 2). For example, consider a scenario where a customer purchases an item. This interaction can be modeled as a heterogeneous graph with the triplet (customer, transaction, item), where a message originating from a customer node is propagated to an item node via a transaction edge in a single hop.
In contrast, relational data graphs are structured differently. While they are heterogeneous and contain multiple node-types, their edge types are not defined based on semantic interactions but rather by primary-foreign key relationships. Consequently, the fundamental structural unit is not necessarily a triplet but a direct pairwise relationship between nodes, represented as (node-type 1, node-type 2). In a relational data graph, the (customer, transaction, item) triplet is modeled by three nodes: a customer node, a transaction node, and an item node, and pairwise interactions between them i.e., (customer, transaction) and (transaction, item). This distinction is critical in both modeling and analysis and requires further investigation.
Our first observation is node-types in relational graphs can be broadly classified into two categories: (i) those with zero or one foreign key and (ii) those with two or more foreign keys. To illustrate, consider the rel-f1 schema, which tracks all-time Formula 1 racing data since 1950 (see Fig. 1). In this schema, Constructor, races, and drivers each have zero foreign keys, while circuits has one. By contrast, constructor_standings, constructor_results, and standings each have two foreign keys, and results and qualifying each have three. Nodes with two or more foreign keys are highlighted in purple. Subgraphs containing only node-types with zero or one foreign key exhibit no special structural patterns and can be treated as general heterogeneous graphs. However, subgraphs that include node-types with multiple foreign keys reveal unique structures, necessitating further analysis.
@@ 3.1 Challenges in Message Passing for Relational Deep Learning
Node-type with multiple (three or more) foreign keys (hub): When node-types have three or more foreign keys, they form star-shaped subgraphs that serve as communication hubs, bridging multiple node-types. For example, as shown in Fig. 1, the results node mediates interactions among constructors-races, races-drivers, and constructors-drivers. These hub nodes inherit the same inefficiencies seen in the two-foreign-key case, where two-hop message passing often leads to redundant information aggregation. Furthermore, the star-shaped connectivity at the schema level induces a hidden second-order clique structure in the data graph—an aspect that standard modeling approaches fail to exploit effectively. Our proposed approach, which is explained next, can leverage these hidden cliques, which form a critical substructure in many high-impact domain graphs. This transition from star-like to clique-like patterns substantially increases connectivity density, reshaping information propagation and complicating message passing dynamics.
3.2 Atomic Routes in Relational Deep Learning

### v2

@@ Abstract
Predictive tasks on relational databases are critical in real-world applications spanning e-commerce, healthcare, and social media. To address these tasks effectively, Relational Deep Learning (RDL) encodes relational data as graphs, enabling Graph Neural Networks (GNNs) to exploit relational structures for improved predictions. However, existing RDL methods often overlook the intrinsic structural properties of the graphs built from relational databases, leading to modeling inefficiencies, particularly in handling many-to-many relationships. Here we introduce RelGNN, a novel GNN framework specifically designed to leverage the unique structural characteristics of the graphs built from relational databases. At the core of our approach is the introduction of atomic routes, which are simple paths that enable direct single-hop interactions between the source and destination nodes. Building upon these atomic routes, RelGNN designs new composite message passing and graph attention mechanisms that reduce redundancy, highlight key signals, and enhance predictive accuracy. RelGNN is evaluated on 30 diverse real-world tasks from RelBench (Fey et al., 2024), and achieves state-of-the-art performance on the vast majority of tasks, with improvements of up to 25%. Code is available at https://github.com/snap-stanford/RelGNN
Machine Learning, ICML
@@ 1 Introduction
Predictive modeling over relational data (multiple tables connected via primary-foreign key relations) is central to numerous real-world applications: e-commerce platforms forecast product demand, music streaming services personalize recommendations, financial institutions assess credit risk, etc. The common strategy to tackle these predictive tasks relies on classical tabular machine learning approaches (Chen & Guestrin, 2016) that often require flattening relational data into a single table through manual feature engineering (Kaggle, 2022). This approach is not only labor-intensive but also leads to a substantial loss of predictive signal, as it oversimplifies the interconnected structure of relational data during the flattening process.
@@
To overcome the limitations of tabular approaches, Fey et al. (2024) introduced Relational Deep Learning (RDL), a new machine learning paradigm that enables end-to-end trainable neural networks to perform predictive modeling directly on relational databases. In RDL, relational data is represented as a graph, where each entity is represented as a node, and the primary-foreign key links between entities define the edges. This graph-based representation allows Graph Neural Networks (GNNs)  (Gilmer et al., 2017; Hamilton et al., 2017) to serve as predictive models, capturing complex relational dependencies that traditional approaches overlook. Complementing this advancement, RelBench (Robinson et al., 2024) provides the first comprehensive benchmark for evaluating and developing RDL models.
Building effective RDL models is essential for tackling predictive tasks, yet remains quite challenging: Relational entity graphs are large-scale, heterogeneous, and temporally dynamic. The models introduced by Robinson et al. (2024) apply standard heterogeneous GNNs (Schlichtkrull et al., 2018; Hu et al., 2020) to relational entity graphs. This design choice can be suboptimal, since standard heterogeneous GNNs are developed for generic multi-relational graphs, where each edge-type denotes a direct semantic interaction between nodes. In relational databases, however, edges are defined by primary–foreign key links that merely record table connectivity, rather than a native semantic relation. Treating these schema-driven links as ordinary semantic edges overlooks the database’s characteristic topology (e.g., many-to-many motifs, as discussed later), misrepresents how information should propagate, and ultimately limits model fidelity. These differences motivate the design of architectures tailored to the structural regularities of relational databases.
@@
Figure 1: An illustration of the key concepts in our method. (a) The primary-foreign key relation of the rel-f1 dataset. Arrows point from a node with a foreign key to the node with the corresponding primary key. Nodes with zero or one foreign key are marked in pink, and the corresponding foreign key is illustrated by a solid line. Nodes with two or more foreign keys are marked in purple, and the corresponding foreign keys are illustrated by dotted lines. (b) An example of the bridge structure, where standings node is a bridge node. (c) An example of the hub structure, where results node is a hub node. (d) Examples of three atomic routes, where the nodes within each box constitute a distinct atomic route.
Our contribution.
@@
To address these challenges, we propose RelGNN, a novel GNN framework that introduces the composite message passing and graph attention mechanism to fully exploit the unique structural properties of graphs built from relational database. RelGNN introduces the concept of atomic routes, which are simple paths that support complete, single-hop information exchange between the source and destination nodes, grounded in primary–foreign key relationships (cf. Figure 1(d)). When a table has a single foreign key, an atomic route consists of a pair of node-types and the edge connecting them. When a table has multiple foreign keys, it induces a set of atomic routes in the form of (source $\rightarrow$ intermediate $\rightarrow$ destination) paths, each of which connects two entities through a shared intermediate node (i.e., a bridge or hub node). Although reminiscent of meta-paths (Sun et al., 2011) used in conventional heterogeneous graphs, atomic routes differ fundamentally in how they are constructed. While meta-paths typically rely on manual design guided by domain expertise, atomic routes are automatically and systematically derived from primary–foreign key relationships in relational databases, making them both broadly applicable and highly scalable across diverse datasets. Building on these routes, RelGNN designs composite message passing with attention mechanism that directly aggregates messages along atomic routes in a single step, avoiding redundant hops and preventing irrelevant information aggregation. This enables more efficient and accurate extraction of predictive signals compared to conventional heterogeneous GNNs.
We assess the performance of the proposed RelGNN across all tasks in RelBench (Fey et al., 2024), a benchmark which spans seven diverse relational databases covering e-commerce, social networks, sports, and medical platforms. RelBench features 30 real-world predictive tasks cast as entity classification, entity regression, and recommendation. RelGNN surpasses all baselines on 27 of the 30 tasks while performing comparably on the remaining three. Notably, RelGNN achieves more than a 4% improvement over a standard heterogeneous GNN in 17 out of 30 tasks, and provides up to a 25% improvement on the site-success regression task in the rel-trial database.
@@ 2.1 Relational Database
A relational database $(\mathcal{T},\mathcal{L})$ consists of a set of tables $\mathcal{T}=\{T_{1},\ldots,T_{n}\}$ and a set of links between them $\mathcal{L}\subseteq\mathcal{T}\times\mathcal{T}$. Each table is a set $T=\{v_{1},...,v_{n_{T}}\}$, where the elements $v_{i}\in T$ are called rows or entities. Each entity $v\in T$ has a unique primary key $p_{v}$ that distinguishes it from other entities within the table. An entity can possess one or more foreign keys. Formally, each foreign key $\mathcal{K}_{v}\subseteq\{p_{v^{\prime}}:v^{\prime}\in T^{\prime}\text{ and }(T,T^{\prime})\in\mathcal{L}\}$ links an entity $v\in T$ to entity $v^{\prime}\in T^{\prime}$, where $p_{v^{\prime}}$ is the primary key of $v^{\prime}$ in table $T^{\prime}$. An entity may have several attributes $x_{v}$, which represent the informational content, and may optionally include a timestamp $t_{v}$, indicating when the corresponding event occurred. Primary keys, foreign keys, attributes, and the timestamp are all realized as table columns. A link $L=(T_{\rm fkey}$, $T_{\rm pkey})$ between tables exists if a foreign key column in $T_{\rm fkey}$ references a primary key column of $T_{\rm pkey}$.
@@
To capture the structural relationships between tables, we introduce the schema graph, which encodes the table-level topology of a relational database. Given a relational database $(\mathcal{T},\mathcal{L})$, the schema graph is defined as $(\mathcal{T},\mathcal{L})$, where each table (i.e., node-type) corresponds to a node, and each primary–foreign key relationship defines a directed edge from the foreign key table to the primary key table. Figure 1 (a) illustrates the schema graph of rel-f1 dataset and Figure 4 shows the schema graphs of all the other datasets in RelBench. This definition corresponds to the directed variant of the schema graph described in  Fey et al. (2024). Note that a node in the schema graph corresponds to a node-type in the relational entity graph, which we introduce in the next subsection.
2.2 Relational Deep Learning (RDL)
Fey et al. (2024) introduced RDL, an end-to-end framework for predictive modeling on relational databases with neural networks. RDL represents the database as a relational entity graph, which is a temporal, heterogeneous graph, where each table becomes a node-type, each row (entity) an individual node, and every primary–foreign key relationship an edge (Figure 2 (b)–(c)). The columns of each table provide the initial feature vector for their corresponding nodes. The transformation from the relational database to its relational entity graph form is lossless, preserving all information contained in the original tables.
Time is a first-class citizen in RDL. Each entity $v$ may carry a timestamp $t_{v}$ that records when the corresponding event occurred—for example, every row in the Transactions table logs the moment of purchase. Many tasks are inherently temporal (e.g., predicting next-week product sales), so the model must respect causality. RDL therefore employs temporal neighbor sampling (Hamilton et al., 2017; Fey et al., 2024): for a seed entity at time $t$, it builds a subgraph that contains only nodes with timestamps $\leq t$, excluding all future information to avoid leakage. A GNN is then trained end-to-end on these time-consistent subgraphs, eliminating the need for manual feature engineering (Figure 2 (d)).
3 Bridge and Hub Topology in RDL (our work)
3.1 Structural Patterns in Relational Databases
Heterogeneous graph topology.
General heterogeneous graphs encode information between nodes and edges of different types. The basic structural unit of the heterogeneous graph schema is a relation, written as a triplet (source node-type, edge-type, destination node-type). For example, a retail purchase is captured by the triplet (customer, transaction, item), which represents a customer purchasing an item. Heterogeneous GNNs (Schlichtkrull et al., 2018; Hu et al., 2020), are designed to process these relations through edge-semantic message passing, e.g., a message originating at a customer node traverses a transaction edge and is aggregated at the corresponding item node in a single hop. This message passing mechanism learns representations of both nodes and edge-types, and preserves the distinct semantics of each relation.
Relational entity graph topology.
Unlike general heterogeneous graphs, where an edge denotes a semantically meaningful relation, the edges of relational entity graphs are formed by references to primary foreign keys in the database schema. Hence, each edge-type carries no intrinsic semantics beyond “table A points to table B,” and the basic structural unit collapses to a pair (node-type 1, node-type 2), rather than a semantic triplet. Consider the retail example: instead of a single relation (customer, transaction, item), the database materializes a junction table transaction. In the relational entity graph, this becomes an explicit intermediate node-type, producing two edges (transaction, customer) and (transaction, item). Because these edges encode only primary-foreign key relationships—not the semantic of purchasing—the usual message passing assumptions for heterogeneous graphs no longer hold. This shift from semantic triplets to purely structural links therefore calls for further investigation.
Many-to-many relationships.
The transition from semantic triplets to primary–foreign key edges significantly impacts how many-to-many relationships are represented in relational databases. A many-to-many relationship describes a scenario where multiple instances of one entity type are associated with multiple instances of another. Although many-to-many relationships are ubiquitous, they cannot be directly encoded using a single primary–foreign key connection, as by definition, a foreign key can reference only one unique primary key. To address this, databases introduce junction tables that decompose many-to-many relationships into pairs of one-to-many links. In the retail example, customers and items form a many-to-many relationship—each customer can purchase multiple items, and each item can be purchased by multiple customers. This association is mediated by a transaction node, which serves as an intermediate connector in the relational entity graph. Each transaction node contains a pair of foreign keys, one referencing a customer node and the other referencing an item node, thereby decomposing the many-to-many association into two one-to-many links.
The previous observations lead to the following categorization of node-types in relational entity graphs: (i) node-types with zero or one foreign key, and (ii) node-types with two or more foreign keys. To illustrate them, consider the rel-f1 schema, which tracks all-time Formula 1 racing data since 1950 (cf. Figure 1 (a)). In this schema, constructors, races, and drivers each have zero foreign keys, and circuits has one. Constructor_standings, constructor_results, and standings each have two foreign keys, and results and qualifying each have three. Node-types with zero or one foreign key, highlighted in pink, do not require intermediate nodes to support message exchange and they exhibit a standard topology. In contrast, node-types with two or more foreign keys, highlighted in purple, serve as structural intermediaries in the schema. Given the ubiquity of many-to-many relationships in real-world interactions, cases involving (ii) are common in relational entity graphs and play a central role in shaping information flow. Next, we analyze the limitations of standard message passing over these structures and motivate a more specialized approach.
3.2 Challenges in Message Passing for RDL
3.2.1 Node-type with two foreign keys (bridge)
[…]
For example, consider the task of predicting race outcomes, where information flows from a races node (source) to a drivers node (destination) via a standings node (bridge) (cf. Figure 1 (b)). In standard two-hop message passing, information from standings reaches the drivers node in both hops, while information from races arrives only in the second. Since the races node typically contains more contextual information relevant to driver performance, this imbalance may degrade performance. As we show in the next section, modeling such structure as an atomic route enables direct one-hop message passing between the relevant node-types without any loss of information.
3.2.2 Node-type with three or more foreign keys (hub)
@@
When a node-type has three or more foreign keys, it forms a star-shaped subgraph that acts as a communication hub, connecting multiple other node-types. For example, as shown in Figure 1 (c), the results node links constructors, races, and drivers, thereby mediating interactions among constructors-races, races-drivers, and constructors-drivers. These hub node-types inherit the same inefficiencies observed in the two-foreign key case (bridge structures): Standard two-hop message passing leads to redundant communication paths and imbalanced predictive signals since information from hub nodes is aggregated multiple times, while more informative signals from source nodes may be underrepresented. Furthermore, the star-shaped connectivity patterns around hub node-types induce latent second-order clique structures at the schema level—structures that are typically overlooked and underutilized by standard modeling approaches. Our proposed approach, detailed next, is designed to exploit these latent clique structures, which constitute critical subgraphs in many real-world domains. This shift from star-shaped to clique-like connectivity substantially increases graph density and transforms the geometry of information flow.
4 Proposed Architecture: RelGNN
4.1 Atomic Routes in RDL